I need help understanding this, Please help. P Varies Directly as R and inversly as the square of S. If P = 175 and R = 252, and S = 6, Find P When R = 196 and S = 7
Answer by scott8148(6628)

you need to find the constant of variation (usually written as k)

P=kR/S^2 __ 175=k*252/36 __ 25=k

P=25*196/49 __ P=100